# Detecting Adversarial Examples

### v1

@@ Abstract
Deep Neural Networks (DNNs) have been shown to be vulnerable to adversarial examples. While numerous successful adversarial attacks have been proposed, defenses against these attacks remain relatively understudied. Existing defense approaches either focus on negating the effects of perturbations caused by the attacks to restore the DNNs’ original predictions or use a secondary model to detect adversarial examples. However, these methods often become ineffective due to the continuous advancements in attack techniques. We propose a novel universal and lightweight method to detect adversarial examples by analyzing the layer outputs of DNNs. Through theoretical justification and extensive experiments, we demonstrate that our detection method is highly effective, compatible with any DNN architecture, and applicable across different domains, such as image, video, and audio.
@@ 1 Introduction
Goodfellow et al. (2014) demonstrated that deep neural networks (DNNs) are vulnerable to adversarial examples and proposed the Fast Gradient Sign Method (FGSM) to craft these adversarial examples by adding perturbations to the model inputs, leveraging the linear nature of DNNs. After the initial introduction of FGSM, various adversarial attacks were proposed across different domains. However, compared to the vast diversity among attack techniques, existing defense methods are built on a few different strategies.
@@ 2.1 Adversarial Attacks
The robustness of deep neural networks and their vulnerability against adversarial examples have been investigated since the introduction of FGSM (Goodfellow et al., 2014). Numerous adversarial attacks have been proposed to generate effective adversarial examples in recent years (Madry et al., 2017; Croce & Hein, 2020; Kurakin et al., 2018; Chen et al., 2017; Ilyas et al., 2018; Mumcu & Yilmaz, 2024a; Wang & He, 2021; Fang et al., 2024; Gao et al., 2020). There are two main adversarial attack settings, namely white-box and black-box. While it is assumed that the attacker has access to the target model in the white-box setting, in the black-box setting, the attacker does not have any prior information about the target model.
White box attacks, including FGSM (Goodfellow et al., 2014), PGD (Madry et al., 2017), APGD (Croce & Hein, 2020), generate adversarial examples by maximizing the target model’s loss function and they are usually referred as gradient-based attacks. BIM (Kurakin et al., 2018) tries to improve gradient-based attack by applying perturbations iteratively. Transferability based black-box attacks, which is introduced in Papernot et al. (2017), are one of the most common black-box approach. The idea is to use an attack on a substitute model for generating adversarial examples for unknown target models, utilizing the transferability of adversarial examples to different DNNs. While adversarial examples generated by these attacks are most effective when the substitute model exactly matches the target model as in a white-box attack setting, their success is shown to be transferable even when there is significant architectural differences between the substitute and target models. Wang & He (2021) introduced VMI and VNI to further extend iterative gradient-based attacks and try to achieve high transferability by considering the gradient variance of the previous iterations. PIF (Gao et al., 2020) uses patch-wise iterations to achieve transferability. ANDA (Fang et al., 2024) aims to achieve strong transferability by avoiding the overfitting of adversarial examples to the substitute model.
@@ 2.2 Adversarial Defenses
Attempting to make changes on the input data for removing the effects of perturbations from adversarial examples is the most common defense strategy. JPEG compression is studied in several works (Cucu et al., 2023; Aydemir et al., 2018; Das et al., 2018), and it is shown that compressing and decompressing helps to remove of adversarial effects on input images. Xie et al. (2017) uses random resizing and padding on the inputs to eliminate the adversarial effects. Xu (2017b) introduces feature squeezing where they use bit reduction, spatial smoothing, and non-local means denoising to detect adversarial examples. Several denoising methods (Liao et al., 2018; Xiong et al., 2022; Salman et al., 2020) were proposed to remove adversarial perturbations from the inputs. Mustafa et al. (2019) uses wavelet denoising and image super resolution as pre-processing steps to create a defense pipeline against adversarial attacks. Prakash et al. (2018) tries to eliminate adversarial effects by redistributing the pixel values via a process called pixel deflection. Adversarial training is another method which is studied to increase robustness of DNNs, however adversarial training often fails to perform well especially under various attack configurations (Bai et al., 2021). In addition, a recent defense strategy which utilizes a baseline model, e.g., Vision Language Models (VLMs), with the assumption that the output of target model and baseline model are close to each other for clean input, but are far away from each other for adversarital input. A recent example is demonstrated by Mumcu & Yilmaz (2024c) where they used CLIP (Radford et al., 2021) to detect adversarial video examples.
@@ 3 Detecting Adversarial Examples
To develop a universal detector that work with any DNN and against any attack, we start with the following generic observation. Although there are numerous attacks with different approaches to generate adversarial examples, all attacks essentially aim to change the model’s prediction by maximizing the loss $\mathcal{L}$ (e.g., cross-entropy loss) between prediction $g(x^{adv})$ and the ground truth $y$ while limiting the perturbation Fang et al. (2024):
@@
Note that $a_{n}(x)=g(x)$ and $a_{n}(x^{adv})=g(x^{adv})$ and $\mathcal{L}(g(x^{adv}),g(x))$ is monotonic with $\|g(x^{adv})-g(x)\|_{\infty}$. Hence, $\|a_{n}(x^{adv})-a_{n}(x)\|_{\infty}$ is maximized while limiting $\|x^{adv}-x\|_{\infty}$ by a small number. Finally, the perturbation aligned with DNN weights is amplified as it sequentially moves through the DNN layers Goodfellow et al. (2014),
@@ 3.1 Layer Regression
While Theorem 1 provides the theoretical motivation, the result in Corollary 1 provides a mechanism to detect adversarial samples if we can train a suitable function $f$ (Figure 1). We make four approximations to obtain a practical algorithm based on Corollary 1. First, we propose to use a multi-layer perceptron (MLP) to approximate $f$. Second, since $a_{n}(x)$ denotes the predicted class probabilities, we choose the feature vector $a_{n-1}$, which takes unconstrained real values, as the target to train MLP as a regression model. Third, we approximate non-differentiable $\|\cdot\|_{\infty}$ with $\|\cdot\|_{2}$ to train the MLP using the differentiable mean squared error (MSE) loss. To empirically check the validity of Theorem 1 under these three approximations, we use the ImageNet validation dataset, ResNet-50 He et al. (2015) as target model, and PGD attack Madry et al. (2017) to compute the mean and standard deviation of normalized change in layer 1 $\tilde{d}_{1}=\frac{\|a_{1}(x^{adv})-a_{1}(x)\|_{2}}{\|a_{1}(x^{adv})\|_{2}+\|a_{1}(x)\|_{2}}$ and layer $n-1$ $\tilde{d}_{n-1}=\frac{\|a_{n-1}(x^{adv})-a_{n-1}(x)\|_{2}}{\|a_{n-1}(x^{adv})\|_{2}+\|a_{n-1}(x)\|_{2}}$.
@@ 4 Experiments and Analyses
Defenses: The following defense methods are used in the experiments: JPEG compression (JPEG) (Das et al., 2018), Randomization (Random)(Xie et al., 2017), Deflection (Deflect) (Prakash et al., 2018), VLAD (Mumcu & Yilmaz, 2024c), Feature Squeezing (FS) Xu (2017a), Wavelet denoising (Denoise) and super resolution (WDSR) (Mustafa et al., 2019). Since VLAD and FS were proposed as detectors in their respective papers, we use their official implementations. The remaining methods were proposed to increase robustness by altering the inputs to remove perturbations. For this reason, we obtain a detection method from them by comparing the predictions after and before the methods are applied. The case the predictions match indicates no attack and a mismatch indicates an attack.
Attacks: BIM (Kurakin et al., 2018), PGD (Madry et al., 2017), PIF (Gao et al., 2020), APGD (Croce & Hein, 2020), ANDA (Fang et al., 2024), VMI and VNI (Wang & He, 2021) attacks are used to generate adversarial examples during the experiments. We consider the most challenging case for the detector, the white-box attack scenario, in which the attacker has full access to the classification model, i.e., no mismatch between the surrogate model and actual model.
Models: To represent the different architectures we used three CNN-based image classification models, VGG19 Simonyan & Zisserman (2014), ResNet50 He et al. (2015), Inceptionv3 Szegedy et al. (2015); and three transformers-based models, ViT Dosovitskiy (2020), DeiT Touvron et al. (2021), LeViT Graham et al. (2021).
Datasets: The experiments are conducted using the validation set of ImageNet dataset (Russakovsky et al., 2015) and the CIFAR-100 dataset (Krizhevsky et al., 2009). For testing, the clean set comprises all images that are correctly classified by the target models. An adversarial set is formed for each attack-target model combination by gathering the attack’s adversarial images misclassified by target model.
@@ 6.1 LR for Video Action Recognition
In this section, we implement LR against video action recognition attacks and compare its performance with existing defenses designed for action recognition models, namely Advit (Xiao et al., 2019), Shuffle (Hwang et al., 2023), and VLAD (Mumcu & Yilmaz, 2024c). In the experiments, PGD-v attack (Mumcu & Yilmaz, 2024c) and Flick attack (Pony et al., 2021) are used to target MVIT (Fan et al., 2021) and X3D (Feichtenhofer, 2020). The experimental settings in Mumcu & Yilmaz (2024c) on Kinetics-400 (Kay et al., 2017) dataset are followed. The results in Table 3 show that LR outperforms the other defenses with an average AUC of 0.93%, followed by VLAD which achieves 0.91%.
@@ 6.2 LR for Speech Recognition
To demonstrate LR’s wide applicability, in addition to computer vision, we also demonstrate its performance against a speech recognition attack. As target model, we use Wav2vec (Schneider et al., 2019) model, which is trained on LibriSpeech (Panayotov et al., 2015) dataset. Model is attacked with FGSM (Goodfellow et al., 2014). LR achieves an average AUC of $0.99$. Figure 7 illustrates some attacked and clean samples from LibriSpeech, along with the MSE values of LR. As shown in the figure, LR can even detect stealthy attacks which do not raise the word error rate (WER) much while distorting the recognized speech. Remarkably, while the WER of the first adversarial example in the figure is lower than the WER of the second clean example ($0.4$ vs. $0.5$), the LR loss for this stealthy adversarial sample is more than ten folds greater than that of the second clean example ($1.6$ vs $0.13$).
@@ Appendix A More Related Works from Other Domains
Adversarial examples are studied on several other domains including video action recognition (Pony et al., 2021), traffic sign recognition (Hsiao et al., 2024), object detection (Zhao et al., 2019), video anomaly detection (Mumcu et al., 2022), and speech recognition (Żelasko et al., 2021). For instance, flickering attack (Flick) (Pony et al., 2021) tries to attack action recognition models by changing the RGB stream of the input videos. Hsiao et al. (2024) investigates the effects of natural light on traffic signs and use it to generate adversarial examples.
Advit (Xiao et al., 2019) is a detection method introduced for videos. It generates pseudo frames using optical flow and evaluates the consistency between the outputs for original inputs and pseudo frames to detect attacks. Another defense method, Shuffle (Hwang et al., 2023), tries to increase the robustness of action recognition models by randomly shuffling the input frames. (Mumcu & Yilmaz, 2024b) proposed to train a lightweight VLM and use it for adversarial traffic sign detection. In speech recognition, defenses like denoising or smoothing are studied against adversarial examples (Żelasko et al., 2021).
@@ Appendix B More details on experimental settings
From the validation set of ImageNet (Russakovsky et al., 2015), we used 40,000 images to train the detectors and the remaining 10,000 for testing. For CIFAR-100 (Krizhevsky et al., 2009) has 100 classes and there are 500 training images and 100 testing images per class, resulting in a total of 50,000 training and 10,000 test images. During the tests, only the correctly classified images by the target models were used. Total number for test images for each specific model and dataset is given in Table 4.
@@
VLAD (Mumcu & Yilmaz, 2024c) was originally proposed for video recognition attacks, with the initial implementation designed for videos consisting of 30 frames. In the experiments conducted in Section 4, we used the official VLAD implementation. However, instead of averaging the scores across 30 frames, we applied it to a single image.
@@ Appendix C Selected subset vectors during LR training
A specific subset of layer vectors $a_{r}$, as described in equation 7, is chosen for each target model that is used during the experiments in Section 4. For the models, Pytorch Image Models (timm) (Wightman, 2019) is used.
@@ Appendix D Pseudo-Code for LR
In section 3.1, we explain our detection algorithm in details. Here, in Algorithm 1, we also provide a PyTorch Paszke et al. (2017) like pseudo code for LR.
@@ Appendix E Visualized adversarial examples
Figure 8: Adversarial examples generated with BIM (Kurakin et al., 2018), PGD (Madry et al., 2017), PIF (Gao et al., 2020), APGD (Croce & Hein, 2020), ANDA (Fang et al., 2024), VMI and VNI (Wang & He, 2021) by attacking target model ResNet50 (He et al., 2015). First column represents the clean samples.
Figure 9: Visualization perturbations generated with BIM (Kurakin et al., 2018), PGD (Madry et al., 2017), PIF (Gao et al., 2020), APGD (Croce & Hein, 2020), ANDA (Fang et al., 2024), VMI and VNI (Wang & He, 2021) by attacking target model ResNet50 (He et al., 2015). First row shows the clean sample and corresponding adversairal example for each attack. Second row demonstrates the noises added to the clean sample by each attack.
Figure 10: Effects of different attack strength values on a sample. The attack strength parameter $\epsilon$ is increased in the direction of the arrow. In the first sample, there is no attack. Samples are generated with adversarial attack PGD (Madry et al., 2017) and target model ResNet50 (He et al., 2015)
@@ Appendix F Additional Ablation study: LR in traffic sign detection
As an additional experiment, we implement LR against attacks that target traffic sign recognition. In this study, ResNet50 is used as target model and attacked with FGSM(Goodfellow et al., 2014), PGD(Madry et al., 2017), Light(Hsiao et al., 2024) and Patch (Ye et al., 2021) attacks. In Table 5, we show that LR achieves an average AUC score of 96%, which further proves the applicability and success of our detection method across different domains.

### v2

@@ Abstract
Deep Neural Networks (DNNs) have been shown to be vulnerable to adversarial examples. While numerous successful adversarial attacks have been proposed, defenses against these attacks remain relatively understudied. Existing defense approaches either focus on negating the effects of perturbations caused by the attacks to restore the DNNs’ original predictions or use a secondary model to detect adversarial examples. However, these methods often become ineffective due to the continuous advancements in attack techniques. We propose a novel universal and lightweight method to detect adversarial examples by analyzing the layer outputs of DNNs. Our method trains a lightweight regression model that predicts deeper-layer features from early-layer features, and uses the prediction error to detect adversarial samples. Through theoretical justification and extensive experiments, we demonstrate that our detection method is highly effective, compatible with any DNN architecture, and applicable across different domains, such as image, video, and audio.
An extended version of this work is published at https://openreview.net/forum?id=0CY5APFnFI. This version is provided for archival purposes only.
@@ 1 Introduction
fgsm demonstrated that deep neural networks (DNNs) are vulnerable to adversarial examples and proposed the Fast Gradient Sign Method (FGSM) to craft these adversarial examples by adding perturbations to the model inputs, leveraging the linear nature of DNNs. After the initial introduction of FGSM, various adversarial attacks were proposed across different domains. However, compared to the vast diversity among attack techniques, existing defense methods are built on a few different strategies.
@@ 2.1 Adversarial Attacks
The robustness of deep neural networks and their vulnerability against adversarial examples have been investigated since the introduction of FGSM (fgsm, ). Numerous adversarial attacks have been proposed to generate effective adversarial examples in recent years (pgd, ; apgd, ; bim, ; chen2017zoo, ; ilyas2018black, ; mumcu2024sequential, ; vmi_vni, ; anda, ; pif, ). There are two main adversarial attack settings, namely white-box and black-box. While it is assumed that the attacker has access to the target model in the white-box setting, in the black-box setting, the attacker does not have any prior information about the target model.
White box attacks, including FGSM (fgsm, ), PGD (pgd, ), APGD (apgd, ), generate adversarial examples by maximizing the target model’s loss function and they are usually referred as gradient-based attacks. BIM (bim, ) tries to improve gradient-based attack by applying perturbations iteratively. Transferability based black-box attacks, which is introduced in papernot2017practical , are one of the most common black-box approach. The idea is to use an attack on a substitute model for generating adversarial examples for unknown target models, utilizing the transferability of adversarial examples to different DNNs. While adversarial examples generated by these attacks are most effective when the substitute model exactly matches the target model as in a white-box attack setting, their success is shown to be transferable even when there is significant architectural differences between the substitute and target models. vmi_vni introduced VMI and VNI to further extend iterative gradient-based attacks and try to achieve high transferability by considering the gradient variance of the previous iterations. PIF (pif, ) uses patch-wise iterations to achieve transferability. ANDA (anda, ) aims to achieve strong transferability by avoiding the overfitting of adversarial examples to the substitute model.
@@ 2.2 Adversarial Defenses
Attempting to make changes on the input data for removing the effects of perturbations from adversarial examples is the most common defense strategy. JPEG compression is studied in several works (jpeg1, ; jpeg2, ; jpeg3, ), and it is shown that compressing and decompressing helps to remove of adversarial effects on input images. random uses random resizing and padding on the inputs to eliminate the adversarial effects. squeezing introduces feature squeezing where they use bit reduction, spatial smoothing, and non-local means denoising to detect adversarial examples. Several denoising methods (denoise1, ; denoise2, ; denoise3, ) were proposed to remove adversarial perturbations from the inputs. wdsr uses wavelet denoising and image super resolution as pre-processing steps to create a defense pipeline against adversarial attacks. deflection tries to eliminate adversarial effects by redistributing the pixel values via a process called pixel deflection. Adversarial training is another method which is studied to increase robustness of DNNs, however adversarial training often fails to perform well especially under various attack configurations (adv_training, ). In addition, a recent defense strategy which utilizes a baseline model, e.g., Vision Language Models (VLMs), with the assumption that the output of target model and baseline model are close to each other for clean input, but are far away from each other for adversarital input. A recent example is demonstrated by vlad where they used CLIP (clip, ) to detect adversarial video examples.
@@ 3 Detecting Adversarial Examples
To develop a universal detector that work with any DNN and against any attack, we start with the following generic observation. Although there are numerous attacks with different approaches to generate adversarial examples, all attacks essentially aim to change the model’s prediction by maximizing the loss $\mathcal{L}$ (e.g., cross-entropy loss) between prediction $g(x^{adv})$ and the ground truth $y$ while limiting the perturbation anda :
@@
Note that $a_{n}(x)=g(x)$ and $a_{n}(x^{adv})=g(x^{adv})$ and $\mathcal{L}(g(x^{adv}),g(x))$ is monotonic with $\|g(x^{adv})-g(x)\|_{\infty}$. Hence, $\|a_{n}(x^{adv})-a_{n}(x)\|_{\infty}$ is maximized while limiting $\|x^{adv}-x\|_{\infty}$ by a small number. Finally, the perturbation aligned with DNN weights is amplified as it sequentially moves through the DNN layers fgsm ,
@@ 3.1 Layer Regression
While Theorem 1 provides the theoretical motivation, the result in Corollary 1 provides a mechanism to detect adversarial samples if we can train a suitable function $f$ (Figure 1). We make four approximations to obtain a practical algorithm based on Corollary 1. First, we propose to use a multi-layer perceptron (MLP) to approximate $f$. Second, since $a_{n}(x)$ denotes the predicted class probabilities, we choose the feature vector $a_{n-1}$, which takes unconstrained real values, as the target to train MLP as a regression model. Third, we approximate non-differentiable $\|\cdot\|_{\infty}$ with $\|\cdot\|_{2}$ to train the MLP using the differentiable mean squared error (MSE) loss. To empirically check the validity of Theorem 1 under these three approximations, we use the ImageNet validation dataset, ResNet-50 resnet as target model, and PGD attack pgd to compute the mean and standard deviation of normalized change in layer 1 $\tilde{d}_{1}=\frac{\|a_{1}(x^{adv})-a_{1}(x)\|_{2}}{\|a_{1}(x^{adv})\|_{2}+\|a_{1}(x)\|_{2}}$ and layer $n-1$ $\tilde{d}_{n-1}=\frac{\|a_{n-1}(x^{adv})-a_{n-1}(x)\|_{2}}{\|a_{n-1}(x^{adv})\|_{2}+\|a_{n-1}(x)\|_{2}}$.
@@ 4 Experiments and Analyses
Defenses: The following defense methods are used in the experiments: JPEG compression (JPEG) (jpeg3, ), Randomization (Random)(random, ), Deflection (Deflect) (deflection, ), VLAD (vlad, ), Feature Squeezing (FS) fs , Wavelet denoising (Denoise) and super resolution (WDSR) (wdsr, ). Since VLAD and FS were proposed as detectors in their respective papers, we use their official implementations. The remaining methods were proposed to increase robustness by altering the inputs to remove perturbations. For this reason, we obtain a detection method from them by comparing the predictions after and before the methods are applied. The case the predictions match indicates no attack and a mismatch indicates an attack.
Attacks: BIM (bim, ), PGD (pgd, ), PIF (pif, ), APGD (apgd, ), ANDA (anda, ), VMI and VNI (vmi_vni, ) attacks are used to generate adversarial examples during the experiments. We consider the most challenging case for the detector, the white-box attack scenario, in which the attacker has full access to the classification model, i.e., no mismatch between the surrogate model and actual model.
Models: To represent the different architectures we used three CNN-based image classification models, VGG19 vgg , ResNet50 resnet , Inceptionv3 inception ; and three transformers-based models, ViT vit , DeiT deit , LeViT levit .
Datasets: The experiments are conducted using the validation set of ImageNet dataset (imagenet, ) and the CIFAR-100 dataset (cifar, ). For testing, the clean set comprises all images that are correctly classified by the target models. An adversarial set is formed for each attack-target model combination by gathering the attack’s adversarial images misclassified by target model.
@@ 6.1 LR for Video Action Recognition
In this section, we implement LR against video action recognition attacks and compare its performance with existing defenses designed for action recognition models, namely Advit (advit, ), Shuffle (shuffle, ), and VLAD (vlad, ). In the experiments, PGD-v attack (vlad, ) and Flick attack (flick, ) are used to target MVIT (mvit, ) and X3D (x3d, ). The experimental settings in vlad on Kinetics-400 (k400, ) dataset are followed. The results in Table 3 show that LR outperforms the other defenses with an average AUC of 0.93%, followed by VLAD which achieves 0.91%.
@@ 6.2 LR for Speech Recognition
To demonstrate LR’s wide applicability, in addition to computer vision, we also demonstrate its performance against a speech recognition attack. As target model, we use Wav2vec (w2v, ) model, which is trained on LibriSpeech (libri, ) dataset. Model is attacked with FGSM (fgsm, ). LR achieves an average AUC of $0.99$. Figure 7 illustrates some attacked and clean samples from LibriSpeech, along with the MSE values of LR. As shown in the figure, LR can even detect stealthy attacks which do not raise the word error rate (WER) much while distorting the recognized speech. Remarkably, while the WER of the first adversarial example in the figure is lower than the WER of the second clean example ($0.4$ vs. $0.5$), the LR loss for this stealthy adversarial sample is more than ten folds greater than that of the second clean example ($1.6$ vs $0.13$).
@@ Appendix A More Related Works from Other Domains
Adversarial examples are studied on several other domains including video action recognition [31], traffic sign recognition [15], object detection [51], video anomaly detection [23], and speech recognition [50]. For instance, flickering attack (Flick) [31] tries to attack action recognition models by changing the RGB stream of the input videos. [15] investigates the effects of natural light on traffic signs and use it to generate adversarial examples.
Advit [44] is a detection method introduced for videos. It generates pseudo frames using optical flow and evaluates the consistency between the outputs for original inputs and pseudo frames to detect attacks. Another defense method, Shuffle [16], tries to increase the robustness of action recognition models by randomly shuffling the input frames. [24] proposed to train a lightweight VLM and use it for adversarial traffic sign detection. In speech recognition, defenses like denoising or smoothing are studied against adversarial examples [50].
@@ Appendix B More details on experimental settings
From the validation set of ImageNet [34], we used 40,000 images to train the detectors and the remaining 10,000 for testing. For CIFAR-100 [19] has 100 classes and there are 500 training images and 100 testing images per class, resulting in a total of 50,000 training and 10,000 test images. During the tests, only the correctly classified images by the target models were used. Total number for test images for each specific model and dataset is given in Table 4.
@@
VLAD [25] was originally proposed for video recognition attacks, with the initial implementation designed for videos consisting of 30 frames. In the experiments conducted in Section 4, we used the official VLAD implementation. However, instead of averaging the scores across 30 frames, we applied it to a single image.
@@ Appendix C Selected subset vectors during LR training
A specific subset of layer vectors $a_{r}$, as described in equation 7, is chosen for each target model that is used during the experiments in Section 4. For the models, Pytorch Image Models (timm) [43] is used.
@@ Appendix D Pseudo-Code for LR
In section 3.1, we explain our detection algorithm in details. Here, in Algorithm 1, we also provide a PyTorch [30] like pseudo code for LR.
@@ Appendix E Visualized adversarial examples
Figure 8: Adversarial examples generated with BIM [20], PGD [22], PIF [11], APGD [4], ANDA [9], VMI and VNI [42] by attacking target model ResNet50 [14]. First column represents the clean samples.
Figure 9: Visualization perturbations generated with BIM [20], PGD [22], PIF [11], APGD [4], ANDA [9], VMI and VNI [42] by attacking target model ResNet50 [14]. First row shows the clean sample and corresponding adversairal example for each attack. Second row demonstrates the noises added to the clean sample by each attack.
Figure 10: Effects of different attack strength values on a sample. The attack strength parameter $\epsilon$ is increased in the direction of the arrow. In the first sample, there is no attack. Samples are generated with adversarial attack PGD [22] and target model ResNet50 [14]
@@ Appendix F Additional Ablation study: LR in traffic sign detection
As an additional experiment, we implement LR against attacks that target traffic sign recognition. In this study, ResNet50 is used as target model and attacked with FGSM[12], PGD[22], Light[15] and Patch [49] attacks. In Table 5, we show that LR achieves an average AUC score of 96%, which further proves the applicability and success of our detection method across different domains.